OpenStudy (anonymous):

Can anyone please explain me the rate equation for a first order reaction....i mean its derivation? I have a very elementary knowledge of calculus but you may use integration and differentiation

OpenStudy (frostbite):

Alright lets assume we have a first order reaction at we set up the following equation: \[v=k[X]\] Now per definition we say, that the rate of a reaction is given by: \[v=-\frac{ d[X] }{ dt }\] Insert that into the equation before and we get the following expression: \[-\frac{ d[X] }{ dt }=k[X]\] This is a fist order differential equation.

OpenStudy (anonymous):

https://www.youtube.com/watch?v=lx3Nlj4ZaJc&NR=1&feature=endscreen this video helped me in some way too....

OpenStudy (anonymous):

but i stil need a complete explanation specific to the question

OpenStudy (frostbite):

The solution to this equation would be: \[\frac{ 1 }{ [X] } d[X]=-k dt\] integrate it over the time t=0 to t=t: \[\int\limits_{[X_{0}]}^{[X]} \frac{ 1 }{ [X] } d[X]=\int\limits_{0}^{t}-k dt\] This is the same as the following when you calculate it: \[\ln([X])-\ln([X_{0}])=-kt\] Rearange it: \[\ln([X])=-kt + \ln([X_{0}])\] Take the exponential function on both sides and we get: \[[X]=[X_{0}]*e ^{-kt}\]